OpenStudy (anonymous):

3. a. Calculate the angular speed of rotation of Earth about its axis (Assume earth to be rigid body) b. If Earth shrinks to half of its radius without change in mass i) Which of the following is conserved? (moment of inertia, angular momentum, angular speed)

OpenStudy (anonymous):

the earth makes one rotation about its axis every 24 hours \[(2*\pi) / 24*60*60= 7.27 * 10^-5 rad/\sec\]... conservation of momentum will be followed...according to me..

OpenStudy (anonymous):

b)angular momentum is conserved

OpenStudy (anonymous):

yeah. angular momentum will be conserved as there is no net torque acting